. An ideal parallel plate capacitor of capacitance C has circular plates located at z = 0 and z = d respectively. The medium between the plates is a linear, homogeneous, isotropic dielectric of dielectric constant K. The capacitor is connected to a resistance R in series, and a voltage V is applied to the circuit. The charge q on the capacitor plates increases with time according to q = C V (1-e−t/RC). Find the magnitude of the
magnetic field H inside the dielectric.
1
Expert's answer
2020-06-29T14:15:38-0400

As per the given question,

Number of turns in the primary coil "(N_1)=500"

Number of turns in the secondary coil "(N_2)=200"

Resistance in the primary coil "(R_1)=0.3\Omega"

Resistance in the secondary coil"(R_2)=0.02\Omega"

Leakage reactance in the primary coil "(X_1)=2\Omega"

Leakage reactance in the secondary coil"(X_2)=0.05\Omega"

Let R_1' be the resistance of the resistance of primary referred to as secondary ,

"\Rightarrow R_1' =R_1(\frac{N_2}{N_1})^2"

"\Rightarrow R_1'=0.3\times (\frac{200}{500})^2=\frac{0.3\times 4}{25}\Omega=0.048\Omega"

Let R_2' be the resistance of the secondary referred to as primary,

"\Rightarrow R_2'=R_2(\frac{N_1}{N_2})^2=0.02\times (\frac{500}{200})^2\Omega"

"=0.02\times 6.28 \Omega =0.1248\Omega"

Let "X_1'" be the leakage reactance of the primary referred to as secondary,

"\Rightarrow X_1'=X_1(\frac{N_2}{N_1})^2=2\times (\frac{200}{500})^2=\frac{8}{25}\Omega=0.32\Omega"

Let"X_2'" be the leakage reactance of the secondary referred to as primary,

"\Rightarrow X_2' =X_2(\frac{N_1}{N_2})^2=0.05\times (\frac{500}{200})^2=0.05\times\frac{25}{4}=0.3125\Omega"

a) Equivalent resistance and reactance referred to as primary,

"R_{eq1}=R_1+R_2'=(0.3+0.1248)\Omega =0.4248\Omega"

"X_{eq1}=X_1+X_2'=2\Omega+0.3125\Omega =2.3125\Omega"

b) Equivalent resistance and reactance referred to as secondary coil,

"R_{eq2}=R_2+R_1'=(0.02+0.048)\Omega =0.068\Omega"

"X_{eq2}=X_2+X_1'=0.05\Omega+0.32\Omega =0.37\Omega"

Equivalent impedance referred to as primary side,

"z=\sqrt{R_{eq1}^2+X_{eq1}^2}=\sqrt{0.4248^2+2.3125^2}=2.35\Omega"